The measure of an angle is 5 times the measure of its supplement. Find the measures of the two angles
Answer by nerdybill(7384)

The measure of an angle is 5 times the measure of its supplement. Find the measures of the two angles
.
Let x = measure of angle
then
180-x = supplement of angle
.
x = 5(180-x)
x = 900-5x
6x = 900
x = 150 degrees (measure of angle)
.
supplement:
180-x = 180-150 = 30 degrees